two angles are supplementary and the larger is 20 degree less than the three times the smaller.Find the angles

Answer by bluemockingjay7(41)

supplementary = 180 degrees
larger = 3x - 20
smaller = x
3x - 20 + x = 180
4x - 20 = 180
4x = 180 + 20
4x = 200
4x/4 = 200/4
x = 50
smaller = 50 degrees
larger = 3(50) - 20 = 150 - 20 = 130 degrees